OpenStudy (anonymous):

Find the solution of the linear equation. 4a - 11 = 2a + 2(a + 6) A. a = 1 B. a = 12/11 C. infinitely many solutions D. no solution @hero

jagr2713 (jagr2713):

No solution

OpenStudy (anonymous):

D

jagr2713 (jagr2713):

Yea

OpenStudy (anonymous):

Don't just give me an answer! One that's against the rules and two I want to KNOW HOW TO DO IT! Can you please shoe me the steps on how you got that answer? @jagr2713 @khurramshahzad

OpenStudy (anonymous):

Any idea how to start it?

jagr2713 (jagr2713):

4a−11=2a+2a+12 Since 2a and 2a are like terms, add 2a to 2a to get 4a. 4a−11=4a+12 Since 4a contains the variable to solve for, move it to the left-hand side of the equation by subtracting 4a from both sides. 4a−11−4a=12 Since 4a and −4a are like terms, add −4a to 4a to get 0. 0−11=12 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. −11=12 Since −11≠12, there are no solutions.

jagr2713 (jagr2713):

HAPPY

OpenStudy (anonymous):

its nothing in it to explain :)

OpenStudy (anonymous):

Can't even teach anymore without someone showing all the steps at once.

jagr2713 (jagr2713):

beat you to the punch lol

jagr2713 (jagr2713):

@CrazyCountryGirl happy!

OpenStudy (anonymous):

Been happening a lot today >_>

jagr2713 (jagr2713):

ikr i hate when that happens

jagr2713 (jagr2713):

i am doing the problem and someone comes and gives the steps faster than me

OpenStudy (anonymous):

Thanks, that's much better. I actually get it now. :)

jagr2713 (jagr2713):

Yw :D